OpenStudy (anonymous):

Answeer For Medal What is the equation of the line that has a slope of 4/5 and a y-intercept of –1?

OpenStudy (anonymous):

The equation for a line is \[y = mx + b\]Where m = slope and b = y intercept. See if you can fill this out if we know the slope and y-intercept! Let me know if you need any more help.

OpenStudy (anonymous):

-4/5+1 is the answer?

OpenStudy (anonymous):

Almost. You got the slope and y intercept right, but the actual equation for the line is y = -4/5x + 1 you just forgot x and y. Does that make sense?

OpenStudy (anonymous):

y=4/5x+1

OpenStudy (anonymous):

Oh yeah, my bad. y = 4/5x + 1 not -4/5x! Nice catch. :)